### v1

Abstract
As large language models (LLMs) become increasingly central to Arabic NLP applications, evaluating their understanding of regional dialects and cultural nuances is essential, particularly in linguistically diverse settings like Saudi Arabia. This paper introduces Absher, a comprehensive benchmark specifically designed to assess LLMs performance across major Saudi dialects. Absher comprises over 18,000 multiple-choice questions spanning six distinct categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition. These questions are derived from a curated dataset of dialectal words, phrases, and proverbs sourced from various regions of Saudi Arabia. We evaluate several state-of-the-art LLMs, including multilingual and Arabic-specific models. We also provide detailed insights into their capabilities and limitations. Our results reveal notable performance gaps, particularly in tasks requiring cultural inference or contextual understanding. Our findings highlight the urgent need for dialect-aware training and culturally aligned evaluation methodologies to improve LLMs performance in real-world Arabic applications.
keywords:
Large Language Models, Natural Language Processing , Dialects Understanding
1 Introduction
Recent breakthroughs in the area of language processing have shown that contemporary computational models can perform a variety of tasks, including text production, question answering, and translation between languages [1, 2]. Most of the models, though, have been developed with a prevailing focus on English and Modern Standard Arabic (MSA), limiting their performance in domains with linguistic and cultural diversity, such as Saudi Arabia. These systems tend to find it difficult to interpret expressions related to local identity like idiomatic expressions, cultural practices, and dialectal tones that are underrepresented in the data used to train them [3, 4].
Saudi Arabia, as depicted in Figure 1, have a unique and rich linguistic landscape characterized by a significant variety of regional dialects. These include Central, Western, Southern, Eastern, and Northern, each with distinct features in pronunciation, vocabulary, and grammar [5]. Beyond phonetic and lexical differences, these dialects are intimately linked to region-specific proverbs, traditional expressions, and patterns of social interaction, forming a fundamental part of ordinary conversation. While essential, these features are rarely incorporated into existing language resources or evaluation tools [6, 7]. Therefore, there is growing academic interest in exploring ways to enable the proper interpretation of such dialectal and cultural features in computational models.
Figure 1: A map of Saudi Arabia showing the different regions and their dialectical variations.
Beyond improving technical accuracy, interpreting Saudi dialects also has broader sociolinguistic significance. These dialects are not only linguistic systems but also carriers of regional identity, social norms, and cultural heritage. Ensuring that LLMs can process and respond to them fairly contributes to linguistic inclusivity and mitigates the marginalization of underrepresented communities. Such efforts are vital for the development of equitable AI systems and for informing culturally aware language technologies and public policy.
Current benchmarks for Arabic language models are based primarily on MSA or generalized applications [8, 9], with very little attention paid to dialect and cultural representation. This under-representation affects the fairness as well as the effectiveness of LLMs when utilized by Arabic speakers in real-world situations. In particular, Saudi dialects and cultural elements are less represented in training sets and test sets, leading to inferior performance on culturally contextualized tasks like the interpretation of idiomatic expressions or answering appropriately in contextually sensitive contexts [10, 11].
To address this critical gap, our work introduces Absher, a novel benchmark specifically designed to assess the performance of large language models (LLMs) on the unique and fine-grained linguistic and cultural nuances of Saudi dialects. As illustrated in Figure 2, Absher provides a comprehensive evaluation framework built upon a curated set of dialectal contents—including words, phrases, and proverbs—representing diverse regions across Saudi Arabia. The benchmark comprises a total of 18,564 multiple-choice questions, systematically generated to cover six distinct types: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
Figure 2: Overview of our benchmark.
In our work, we aim to provide answers to these key questions. First, we investigate the ability of LLMs to interpret Saudi dialects and cultural expressions by creating a new benchmark dataset that captures the nuances of both. Second, we identify the main challenges LLMs face in this context through a detailed error analysis of their performance on this benchmark. Finally, we present a comparative analysis of different LLMs, evaluating their strengths and weaknesses in handling Saudi dialect and culture-specific tasks.
Our work addresses the aforementioned research questions through the following core objectives:
  • Benchmark Development: Construct a culturally grounded benchmark that systematically evaluates LLMs across diverse Saudi dialects. The benchmark includes multiple question formats targeting both linguistic comprehension and cultural awareness, enabling fine-grained assessment.
  • Performance Evaluation: Conduct a rigorous evaluation of several LLMs using the constructed benchmark to assess their capabilities in interpreting Saudi dialects. The analysis identifies their strengths, limitations, and areas that require further research and refinement.
2 Related Work
2.1 Arabic Dialects in LLMs
Recent advances in Arabic NLP have led to increased attention toward evaluating large language models (LLMs) on dialectal and cultural understanding. Several benchmarks have been proposed to assess model performance in these dimensions. For instance, AraDiCE [7] introduced a benchmark targeting dialectal and cultural understanding across Levantine, Egyptian, and Gulf Arabic, revealing critical limitations in generation, translation, and reasoning capabilities. Calibration-Arabic-PLMs [12] focused on model confidence calibration in dialect identification tasks, highlighting inconsistencies between model certainty and actual performance. Health-LLM-Arabic [13] evaluated LLMs like GPT-4 on health-related claims involving Saudi, Egyptian, and Moroccan dialects, and proposed the Cultural Sensitivity Score to quantify dialect adaptation in clinical scenarios.
Complementing these evaluation-focused efforts, a number of datasets and models have been introduced to enhance dialect and culture awareness in Arabic LLMs. Dallah [14] presented a dialect-aware multimodal assistant fine-tuned across six Arabic dialects. AraReasoner [15] provided a benchmarking suite for assessing the reasoning capabilities of LLMs across a wide range of Arabic NLP tasks. Jawaher [16] curated a multidialectal dataset of Arabic proverbs designed to test LLMs’ understanding of figurative and culturally grounded expressions. Similarly, CARE [17] introduced a multilingual benchmark with human preference ratings aimed at evaluating cultural alignment in Arabic and Chinese contexts. Palm [18] contributed a linguistically and culturally diverse dataset spanning 22 Arab countries, with instructions formulated in both MSA and regional dialects to support comprehensive evaluation.
In the multimodal space, several resources have emerged to test vision-language capabilities in culturally rich Arabic settings. Peacock [19] introduced a family of Arabic multimodal LLMs and released Henna, a benchmark designed to assess cultural competence. JEEM [20] focused on visual question answering and image captioning across four Arabic dialects, using culturally specific imagery. Lastly, Swan and ArabicMTEB [21] offered a family of Arabic-centric embedding models and a robust benchmark suite for evaluating cross-lingual, multi-dialectal, and multi-domain embedding quality.
2.2 Saudi Dialects in LLMs
Despite growing interest in Arabic NLP, few studies have explicitly focused on evaluating large language models (LLMs) within the context of Saudi dialects. One of the first targeted efforts is SaudiCulture [6], a benchmark designed to assess LLMs’ cultural competence across five major regions of Saudi Arabia. Evaluations of models such as GPT-4, LLaMA 3.3, FANAR, Jais, and AceGPT revealed notable performance degradation when addressing region-specific or culturally grounded questions.
Similarly, AL-QASIDA [22] evaluated nine LLMs across eight Arabic dialects, including Saudi Arabic, analyzing generation quality through fidelity, fluency, and diglossia metrics. The study highlighted the impact of pre-training data and suggested methods for dialect-aware model adaptation. In broader benchmarks such as Dallah [14] and Palm [18], Saudi dialects were included alongside other regional variants, though without a focused or fine-grained analysis of Saudi-specific linguistic features.
Beyond direct evaluation, other efforts have explored data creation and translation to support dialectal Arabic processing. For instance, LLM-MT-Data [23] employed LLMs to translate dialectal Arabic—sourced from the Saudi Audio Dataset for Arabic (SADA)—into Modern Standard Arabic, demonstrating the potential of LLMs in generating high-quality parallel data. Similarly, AraBench [24] tackled dialect-specific machine translation using fine-tuning and back-translation techniques to improve Arabic–English translation quality.
While these initiatives mark important progress, most prior work has either concentrated on broader Arabic dialects or MSA, without a dedicated focus on the linguistic and cultural intricacies of Saudi dialects. Dialects such as Central, Western, and Southern are deeply intertwined with region-specific expressions, proverbs, and cultural practices that remain underexplored in current benchmarks. Notably, no prior study has systematically examined Saudi dialects across diverse linguistic units—such as words, phrases, and proverbs—or assessed LLMs’ ability to interpret them in context. This gap underscores the need for a benchmark like Absher, which provides a structured, culturally grounded evaluation of LLMs within the Saudi dialect landscape.
3 Constructing Absher Benchmark
To assess the ability of computational systems to understand the dialectal and cultural nuances specific to Saudi Arabia, we introduce the Absher Benchmark. This evaluation framework is built upon a curated set of dialectal contents—including words, phrases, and proverbs—representing diverse regions across Saudi Arabia. The development process followed a structured pipeline encompassing data collection, task formulation, prompt engineering, automated question generation, and expert-based review as seen in Figure  3.
Figure 3: The overall pipeline of constructing Absher benchmark.
The resulting benchmark provides a systematic and scalable method for evaluating LLMs on their ability to interpret regional language variation, contextual meanings, and culturally embedded expressions. As illustrated in Figure 3, the development process spans from initial data acquisition to final model evaluation, allowing both fine-grained analysis and performance benchmarking of LLMs in Saudi dialects.
3.1 Data Collection
To evaluate the ability of computational models to interpret Saudi dialects and culturally embedded expressions, we introduce Absher, a benchmark grounded in authentic regional content. It is built from a curated collection of dialectal items—including words, phrases, and proverbs—representing linguistic diversity across Saudi Arabia. The benchmark construction followed a structured pipeline encompassing data collection, content classification, task formulation, prompt engineering, automated question generation, and expert validation, as illustrated in Figure 3.
The core of the dataset was collected from the Moajam website [^1] https://ar.mo3jam.com/, a digital repository that aggregates dialectal vocabulary across Arabic regions. We scraped and organized entries into three categories: words, phrases, and proverbs, each annotated with its meaning, dialect label, contextual usage, and translation. This process ensured comprehensive coverage of both literal and figurative expressions relevant to regional Saudi Arabic.
Words refer to vocabulary items specific to Saudi dialects, often including slang or regional variants not found in Modern Standard Arabic—for example, ”Rabshah”.
Phrases encompass idiomatic or conversational expressions commonly used across regions, such as ”Allah Yta’anni”.
Proverbs are culturally rooted sayings that convey implicit meanings, requiring contextual and cultural understanding—e.g., ”Ally ma ya’arf alsger yshwaeh”.
The raw collection initially consisted of 5,483 items, which were refined through preprocessing and filtering, resulting in 3,094 unique contents used to generate over 18,000 multiple-choice questions across six task types. Table 1 summarizes the dataset statistics.
Category # Contents (Raw) # Contents (Filtered) # Generated Qs
Words 4,428 2,533 15,198
Phrases 869 478 2,868
Proverbs 186 83 498
Total 5,483 3,094 18,564
Table 1: Content statistics before and after preprocessing and the number of generated questions by category.
3.2 Pre-processing Steps
After data collection, we applied a systematic pre-processing pipeline to refine the dataset for effective question generation. This phase included deduplication, normalization, and structural formatting. We focused exclusively on entries labeled as Saudi dialects, retaining regional variants such as Central, Western, Southern, Northern, and Eastern. Additionally, general Saudi terms—expressions widely used across the country regardless of region—were preserved to ensure comprehensive linguistic coverage. We further standardized spelling and usage variations across dialects to ensure consistency, particularly for words with multiple colloquial forms. This harmonization was essential for generating uniform prompts and minimizing ambiguity in model evaluation. After filtering and normalization, the dataset was reduced to 3,094 contents: 2,533 words, 478 phrases, and 83 proverbs, as shown in Table 1. Finally, the cleaned data was structured into a machine-readable format suitable for automated prompt generation, forming a high-quality foundation for the benchmark tasks.
3.3 Prompt Design
To rigorously assess the capabilities of LLMs in interpreting Saudi dialects, we employed a structured prompt design strategy tailored to dialectal understanding. Each prompt template included four core components: a task description, a dialect-based question, multiple-choice answer options, and a designated correct answer (see Table 2). Prompts were derived from dialectal contents—words, phrases, and proverbs—each annotated with its dialect label, meaning, and a contextual usage example.
The initial prompt generation followed a simple formulation process, but several issues emerged. Correct answers frequently appeared in fixed positions (e.g., always option A), True/False questions were heavily biased toward “True,” and some questions included incorrect dialectal labels. These inconsistencies introduced unintended biases, undermining the fairness and reliability of the benchmark.
To address these limitations, we refined the prompt design by introducing structural controls. Correct answer positions were randomized to prevent positional bias, and the distribution of True/False items was balanced to ensure fairness. Dialect annotations were reviewed and validated by linguistic experts to confirm their accuracy. These refinements significantly improved the robustness, validity, and representativeness of the final set of multiple-choice questions, establishing Absher as a dependable tool for dialect-sensitive LLMs evaluation. For the complete prompt template, please refer to Appendix A.
Component Description
Task: Describes the task or the function of the question.
Question: The question that is presented to the model.
Multi-choice Options: Option A, B, C, and D.
True/False Options: A) True, B) False
Correct Answer: The correct answer to the question.
Table 2: Template for Question Formulation
3.4 Question Generation
Building upon the refined prompt templates, we conducted large-scale question generation using the ChatGPT API (GPT-4o). We selected GPT-4o for this task due to its strong multilingual capabilities and demonstrated proficiency in both literal and contextual understanding. Its robustness in handling diverse linguistic inputs made it particularly well-suited for generating high-quality, dialect-specific multiple-choice questions that require nuanced cultural reasoning. Compared to smaller or monolingual models, GPT-4o also exhibited greater consistency and fewer hallucinations—qualities essential for ensuring the validity of benchmark items.
Each cleaned and preprocessed content item—whether a word, phrase, or proverb—was programmatically paired with its dialect label, meaning, and usage example to construct a complete input prompt. This setup enabled GPT-4o to generate questions that evaluated not only surface-level understanding but also deeper comprehension of dialectal and cultural nuances. For each dialectal item, the model generated six distinct types of questions, with each type targeting a specific dimension of language understanding:
  • Meaning: Assessing the model’s ability to correctly interpret or define dialectal expressions.
  • True/False: Testing the factual correctness of statements about the term or its usage.
  • Fill-in-the-Blank: Evaluating the model’s ability to infer missing words in context.
  • Contextual Usage: Determining whether the model can accurately apply the term within a complete sentence.
  • Cultural Interpretation: Probing the understanding of cultural or metaphorical meanings, particularly in proverbs.
  • Location Recognition: Identifying the geographic or regional origin associated with the dialectal term.
This systematic generation approach ensured comprehensive coverage of the linguistic and cultural features embedded in Saudi dialects. By applying the process across all content categories—words, phrases, and proverbs—a total of 18,564 multiple-choice questions were generated. These questions form the core of the Absher benchmark and enable a rigorous, multifaceted evaluation of LLM performance. Table 3 shows the distribution of questions across regional dialects, highlighting the benchmark’s broad linguistic representation.
Dialect # Qs
Center 5,652
Western 5,262
Southern 2,682
Northern 414
Eastern 216
General Saudi Terms 4,338
Table 3: Number of questions per regional dialects in the Absher benchmark.
Figure 4 illustrates the six question types included in the benchmark. Together, these types enable a well-rounded evaluation of model capabilities, covering meaning interpretation, cultural and regional understanding, and recognition of local expressions. The use of a standardized prompt template throughout the generation process ensures consistency, minimizes bias, and strengthens the reliability of the evaluation.
Figure 4: An illustration of the different question types. The first line in each example indicates the task type, the second line presents the question, and the subsequent lines list the answer options. The correct answers are marked in green.
3.5 Human Evaluation
To validate the reliability, correctness, and cultural appropriateness of the questions in the Absher benchmark, we conducted a structured human evaluation. This step was essential to ensure that the automatically generated content adhered to linguistic standards and dialectal relevance, especially given the complexity and regional variability of Saudi dialects. Four native Saudi Arabic speakers participated in the evaluation. All annotators were fluent in multiple regional dialects and possessed strong familiarity with the cultural and linguistic diversity across Saudi Arabia. They were divided into two independent groups, each assigned a distinct subset of the data. This setup enabled broader dataset coverage, minimized cross-group bias, and reduced cognitive fatigue.
A pilot study preceded the main annotation phase to ensure consistent interpretation of the evaluation guidelines. Annotators were presented with 20 representative sample contents (120 questions total), covering diverse dialects and question types. These examples served to familiarize them with the annotation tasks and resolve any ambiguities in advance. To test attentiveness and accuracy, the pilot set included quality control questions with plausible yet incorrect answers. Annotators who misclassified these were given follow-up clarifications to align their understanding with the evaluation standards. This step significantly improved consistency and judgment reliability in the formal evaluation.
The main evaluation involved a stratified random sample comprising 10% of the benchmark contents—Words, Phrases, and Proverbs—across all six question types. This resulted in 600 annotated questions. Complete annotation guidelines, including task definitions, assessment criteria, and illustrative examples, are provided in Appendix B. Based on these instructions, annotators completed two core tasks:
  • Question Quality: Evaluate each question for clarity, fluency, and logical structure, with attention to dialectal appropriateness and cultural sensitivity.
  • Answer Correctness: Verify whether the answer labeled as correct by the model aligns with the ground-truth answer based on the dialectal dataset.
Inter-annotator agreement was measured using Cohen’s Kappa coefficient. Table 4 shows strong agreement across most question types, particularly for Meaning and Cultural Interpretation, where Kappa values exceed 0.94. While Fill-in-the-Blank and Contextual Usage tasks yielded slightly lower agreement in question quality (0.71 and 0.72, respectively), the values still fall within acceptable reliability thresholds. Additionally, we computed the overall alignment between model-generated answers and human judgments, which reached 93.03%, confirming the reliability and consistency of the benchmark’s human evaluation process.
Evaluation Meaning True/False FillB Cont.Usage Cult.Interpretation Loca. Recognition
Q-Quality 0.97 0.95 0.71 0.72 0.83 0.85
A-Correctness 0.97 0.87 0.91 0.88 0.94 0.92
Table 4: Cohen’s Kappa scores across the two evaluated criteria. FillB, Cont., Cult., and Loca. refer to “contextual”, “cultural”, “location”, and “Fill-in-the-Blank”.
4 Experiments
We evaluated a set of open-source LLMs: LLaMA-3 8B Instruct [25], Jais-13B [26], ALLaM-7B [27], Mistral-7B [28], Qwen2.5-7B Instruct [29], and AceGPT-7B-chat [30]. These models were chosen based on three main criteria: open availability, Arabic language support, and competitive performance across a range of NLP tasks. All models were evaluated in a zero-shot setting, without any task-specific fine-tuning, to assess their inherent ability to understand Saudi dialects and cultural content. Since the benchmark is composed exclusively of multiple-choice questions, we employed four standard classification metrics: Accuracy, Precision, Recall, and F1 Score. These metrics offer a comprehensive and balanced evaluation of model performance, in line with prior work on zero-shot LLMs assessment for low-resource languages [31].
Model Content Type Accuracy Precision Recall F1-score
[Uncaptioned image] ALLaM Word 40.32% 33.07% 26.63% 17.28%
Phrase 42.28% 30.54% 26.38% 19.42%
Proverb 48.74% 39.95% 37.76% 31.43%
[Uncaptioned image] LLaMA Word 41.30% 36.19% 23.99% 16.63%
Phrase 43.41% 29.66% 24.13% 18.07%
Proverb 45.31% 28.55% 28.60% 18.87%
[Uncaptioned image] Jais Word 42.66% 32.78% 24.76% 14.78%
Phrase 41.94% 37.81% 27.27% 16.27%
Proverb 41.81% 17.52% 28.33% 15.91%
[Uncaptioned image] Mistral Word 47.90% 22.80% 30.05% 23.98%
Phrase 44.80% 28.11% 33.09% 26.02%
Proverb 45.10% 18.27% 29.05% 22.31%
[Uncaptioned image] Qwen Word 63.00% 42.07% 46.10% 41.18%
Phrase 43.68% 33.58% 32.93% 26.00%
Proverb 44.38% 31.36% 33.16% 29.56%
[Uncaptioned image] ACEGPT Word 41.90% 33.98% 33.18% 28.30%
Phrase 39.93% 33.68% 32.55% 29.61%
Proverb 34.81% 37.93% 33.15% 27.51%
Table 5: Average Evaluation Results of Models Performance
4.1 Results
Table 5 presents the performance of the six evaluated models—ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT—across different content types, namely words, phrases, and proverbs. The results reveal noticeable variations in performance depending on both the linguistic content and the complexity of the task [^2] The full set of results is provided in Appendix C..
For word-level content, the Qwen model demonstrated the strongest capabilities, achieving an accuracy of 63.00% and an F1-score of 41.18%. These results indicate Qwen’s effectiveness in handling isolated lexical items, suggesting its advantage in processing tasks that rely on clear, context-independent meanings. In contrast, phrase-level tasks presented a more competitive landscape. Mistral achieved the highest F1-score at 26.02%, with Qwen and ACEGPT following closely behind. This reflects a broader challenge for all models when dealing with phrases, which often carry more ambiguous or context-dependent meanings compared to individual words. Proverb-based tasks emerged as the most challenging across all models, reflecting the complexity and cultural depth embedded in such expressions. Despite these challenges, ALLaM led in this category with an accuracy of 48.74% and an F1-score of 31.43%, indicating its relative strength in handling idiomatic and culturally grounded language. In contrast, models such as Jais consistently recorded lower F1-scores across all content types, particularly for proverbs, highlighting limitations in their ability to process culturally rich or metaphorical expressions.
These findings show a consistent trend: while certain models demonstrate strong performance on words or phrases, accurately processing complex cultural constructs like proverbs remains a significant challenge. This highlights the need for further research to enhance LLMs’ ability to handle the nuanced and context-heavy aspects of language, particularly those rooted in cultural knowledge and idiomatic usage.
5 Analysis
5.1 Impact of Model Size and Language Orientation
Table 6 summarizes the models evaluated with respect to parameter size, language orientation, and open-source availability. The set of models includes Arabic-specific and multilingual LLMs, spanning different sizes and scales, from 7B to 13B parameters.
Model Size Language Accuracy [Uncaptioned image] Qwen2.5 Instruct 1st 7B Multilingual 50.35% [Uncaptioned image] Mistral 2nd 7B Multilingual 45.93% [Uncaptioned image] ALLaM 3rd 7B Arabic 43.78% [Uncaptioned image] LLaMA-3 Instruct 8B Multilingual 43.34% [Uncaptioned image] Jais 13B Arabic 42.14% [Uncaptioned image] AceGPT-chat 7B Arabic 38.88%
Table 6: Models Overview, and Average accuracy
While it is often assumed that larger parameter counts lead to better performance, our results demonstrate that accuracy does not scale linearly with size. For example, Qwen, a 7B multilingual model, achieved the highest overall accuracy of 50.35%, outperforming the larger Arabic-native Jais model (13B) by nearly 10%. Similarly, Mistral (7B) achieved 45.93%, exceeding the performance of some larger Arabic-specialized models. A deeper comparison of Arabic-native models (ALLaM, Jais, ACEGPT) versus multilingual LLMs (Qwen, Mistral, LLaMA) reveals additional insights. Despite being trained specifically on Arabic data, Arabic-native models did not consistently outperform their multilingual counterparts in tasks involving dialectal variation and cultural nuances.
Qwen demonstrated strong performance across tasks, particularly in contextual usage (64.7%) and cultural interpretation (63.7%), highlighting the benefits of multilingual pretraining combined with advanced architecture. ALLaM, though smaller and Arabic-focused, performed competitively in proverb-based cultural interpretation (60.0%) and contextual usage tasks. Mistral, a multilingual model with only 7B parameters, achieved balanced performance and notably excelled in binary reasoning (58.4% in word-level True/False questions), emphasizing the importance of architectural optimization. LLaMA showed clear strengths in location recognition (66.3%) but underperformed in tasks requiring deeper semantic understanding. Among Arabic-native models, ACEGPT achieved solid results in contextual usage (54.5%), while Jais led in location recognition (67.6%) but exhibited limitations in cultural and semantic tasks.
Overall, these findings suggest that specialization in Arabic does not inherently guarantee superior performance on culturally complex or dialectally rich tasks. Instead, the combination of diverse multilingual training, effective architectural design, and optimized parameter utilization appears to be more influential. Performance tends to be stronger on lexical and recognition tasks but decreases as linguistic and cultural complexity increases, particularly with proverbs and culturally loaded expressions.
5.2 Examination of Model Responses by Region
Our evaluation of Large Language Models (LLMs) responses to region-specific linguistic questions, detailed in Figure 5, reveals considerable performance variability influenced by both the geographic origin of expressions and the nature of the linguistic task. Notably, we observed strong performance in the General category for a widely known proverb, with five out of six models providing correct answers. This aligns with the expectation that frequently cited expressions, which appear widely in public discourse, are more likely to be accurately interpreted due to their higher representation in pretraining datasets.
Figure 5: Model responses to dialect-based questions from different Saudi regions. Each example shows the question, its regional, and the answers from six LLMs.
In contrast, expressions from the Southern region resulted in a consistent failure across all six models. The proverb tested was regionally specific and less documented, suggesting limited representation in training data and highlighting the persistent challenges LLMs face with culturally embedded content from low-resource dialects. For the Central region, performance on a location recognition task was mixed. While ALLaM and Jais correctly identified the dialectal origin, other models defaulted to major urban centers such as Riyadh or Jeddah. This pattern likely reflects an urban-centric bias within the models’ training corpora. Interestingly, the Northern region, despite being underrepresented in online resources, saw three models correctly answer a true/false phrase question. This suggests that task structure can, to some extent, mitigate data scarcity, as binary questions allow models to leverage contextual inference or statistical patterns to produce correct responses.
In the Eastern region, performance remained low, with only ACEGPT correctly interpreting the contextual usage of a time-related word. This outcome emphasizes the difficulty models encounter when handling dialectal expressions that rely on contextual understanding, particularly for less represented regions. Finally, for the Western region, five of six models successfully identified the cultural function of a widely circulated annoyance phrase. This likely reflects the increased presence of such expressions in social or digital spaces, enhancing their visibility within pre-training data.
Overall, these results reveal persistent challenges in capturing region-specific linguistic and cultural nuances. Although models tend to perform well with broadly known or globally circulated expressions, they consistently struggle with dialects and phrases deeply rooted in local practices. This underscores the need for more inclusive and dialect-rich datasets that reflect the full diversity of linguistic and cultural realities between regions.
5.3 Performance Across Content and Question Types
Figure 6 presents a detailed analysis of model performance across content categories, revealing key trends in how different LLMs handle dialectal material. Qwen consistently outperformed other models, achieving the highest accuracy on word-level questions, particularly in the meaning and contextual usage types, with an overall peak of approximately 63.5%. ALLaM demonstrated notable strengths in cultural interpretation tasks, especially when dealing with proverbs, where it surpassed 48% accuracy. An interesting pattern emerged in relation to content type: questions based on proverbs consistently resulted in higher model accuracy compared to those based on isolated words or phrases. This may be attributed to the richer context and more explicit cultural cues embedded within proverbs, which provide clearer semantic and pragmatic signals.These results suggest that LLMs benefit from contextually grounded input and perform more reliably when exposed to culturally meaningful expressions.
Complementing this content-based analysis, Figure 7 illustrates how the models performed on different types of questions. Among all types, location recognition questions were the most successfully handled, with average model accuracy exceeding 58%. Jais performed particularly well in this category, reaching 62% accuracy—indicating that models are generally capable of linking dialectal features to their corresponding regions. In contrast, the True/False questions pose the greatest challenge, especially when derived from short or context-poor input, such as individual words or brief phrases. These questions required binary reasoning with limited context, leading to lower performance. Accuracy for this type ranged from 22% (Jais) to 58% (Qwen), underscoring the difficulties LLMs face in making confident decisions under ambiguity or minimal input cues. Overall, these results highlight the dual importance of linguistic structure and contextual richness in benchmarking the dialectal and cultural understanding of LLMs.
Figure 6: Average Accuracy for Different Models by Content Type
Figure 7: Average Accuracy for Different Models by Question Type
Content Type: Proverb
Question Type: Meaning
Region: General
Question Example {RLtext} المهمة: اختر إجابة واحدة فقط دون أي ملاحظات إضافية.
السؤال: ماذا يعني المثل إذا حجت البقرة على قرونها؟
الخيارات:
أ) شيء يمكن تحقيقه إذا تمت الظروف بشكل جيد
ب) شيء مستحيل أو غير قابل للتحقيق
ج) شيء يحدث في حالة نادرة
د) شيء يجب أن يتم في الوقت المحدد
الإجابة الصحيحة: ب
Models’ Answer: [Uncaptioned image] ALLaM: B , [Uncaptioned image] LLaMA: B, [Uncaptioned image] Jais: B, [Uncaptioned image] Mistral: C, [Uncaptioned image] Qwen: B, [Uncaptioned image] ACEGPT: B
Table 7: Model answers to a proverb meaning question in the general dialect.
Content Type: Proverb
Question Type: Fill-in-the-Blank
Region: Southern
Question Example {RLtext} المهمة: اختر إجابة واحدة فقط دون أي ملاحظات إضافية.
السؤال: خطط في _ _ _ واقبص في حيد.
الخيارات:
أ) التراب
ب) الورق
ج) الماء
د) الصخر
الإجابة الصحيحة: ج
Models’ Answer: [Uncaptioned image] ALLaM: D , [Uncaptioned image] LLaMA: B, [Uncaptioned image] Jais: D, [Uncaptioned image] Mistral: D, [Uncaptioned image] Qwen: A, [Uncaptioned image] ACEGPT: D
Table 8: Model answers to a fill-in-the-blank proverb in the southern dialect.
Content Type: Word
Question Type: Location Recognition
Region: Central
Question Example {RLtext} المهمة: اختر إجابة واحدة فقط دون أي ملاحظات إضافية.
السؤال: أين يُستخدم المصطلح أويلاوه بشكل أكثر شيوعاً؟
الخيارات:
أ) القصيم
ب) الرياض
ج) جدة
د) الدمام
الإجابة الصحيحة: أ
Models’ Answer: [Uncaptioned image] ALLaM : A , [Uncaptioned image] LLaMA : B, [Uncaptioned image] Jais : A, [Uncaptioned image] Mistral : B, [Uncaptioned image] Qwen : C, [Uncaptioned image] ACEGPT : D
Table 9: Model answers to a Location Recognition Word question in the Central dialect.
5.4 Case Study of LLMs Responses Across Question Types and Dialects
Our case study investigates the performance of six LLMs across culturally embedded tasks derived from various Saudi dialectal regions. To illustrate observed performance patterns, we examine specific model responses across varied content types, question structures, and dialects, as detailed in Tables 7 to 9. The results indicate that regionally common and widely used expressions are more accurately interpreted by LLMs. For example, in the general category, the proverb:
{RLtext}
“إذا حجت البقرة على قرونها”
“If the cow performs Hajj on its horns” (Table 7) was correctly understood by five out of six models: ALLaM, LLaMA, Jais, Qwen, and ACEGPT all provided the correct answer, while only Mistral failed. This proverb is relatively widespread, and its meaning—“something impossible to achieve”—is often documented in public sources, likely increasing its representation in model training data. In contrast, proverbs that are more regionally isolated or culturally specific, such as the Southern fill-in-the-blank proverb :
{RLtext}
“خطط في ماء واقبص في حيد”
“Plan in water and pinch in a rock” (Table 8), saw a complete failure across all six models. None of the models—ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT selected the correct option, suggesting a consistent gap in exposure to underrepresented regional idioms, especially those originating from areas with fewer accessible textual resources.
Additionally, the study revealed notable differences in the models’ ability to comprehend regional dialects, particularly in how they responded to specific tasks based on region and question type. For example, when tasked with identifying a location-specific term (Table 9) from the Central region, only two models, ALLaM and Jais, were able to correctly associate it with the Al-Qassim region. The other models, however, defaulted to more familiar urban areas, like Riyadh and Jeddah, likely due to the prevalence of these dialects in the training materials. Urban dialects from cities such as Riyadh, Jeddah, and Khobar are more commonly represented in online resources, which could explain why the models showed a preference for these regions. In contrast, like Qassim, Aseer, or the southern regions face challenges due to limited textual data from these areas.
Moreover, when it comes to other local dialects from the Eastern, Northern, or Southern parts of Saudi Arabia, significant discrepancies in the models’ performance were noted. Expressions and words specific to these regions were often misunderstood, highlighting the gap in representation for these dialects. This contrast underscores the fact that while widely recognized expressions are easily interpreted, culturally unique idioms that are specific to certain regions are much harder for the models to grasp.
Overall, our comprehensive evaluation underscores that while contemporary LLMs possess considerable linguistic breadth in Arabic dialects, their depth of understanding for more nuanced, regionally isolated, or culturally embedded expressions remains a significant limitation. These findings collectively point to the crucial need for more extensive and geographically diverse dialectal data in future LLMs training to achieve truly robust and culturally sensitive linguistic AI.
6 Discussion
This research presents Absher, a novel benchmark aimed at evaluating LLMs on Saudi dialects—a linguistically rich yet underrepresented subset of Arabic. Unlike prior work such as Dallah [14], Palm [18], and AL-QASIDA [22], which aggregate dialects or focus on broader Arabic settings, Absher uniquely provides fine-grained, region-specific coverage of Saudi dialects, including Central, Weastern, Southern, Eastern, and Northern varieties.
Our findings highlight notable patterns in how LLMs handle Saudi dialects. Location Recognition tasks yielded the highest performance, likely due to the concrete and frequently occurring nature of geographic references in training data, which aids retrieval and identification. In contrast, True/False questions were the most challenging—especially in low-context settings—reflecting the inherent difficulty of binary reasoning tasks that require nuanced cultural and contextual understanding beyond surface-level lexical cues.
Furthermore, Qwen’s strong performance—especially in Meaning and Contextual Usage questions—demonstrates its robust semantic understanding and ability to accurately interpret dialectal expressions in context. Across content types, tasks involving Proverbs consistently achieved higher accuracy than those based on individual Words or Phrases. This suggests that models benefit from the richer contextual and cultural cues provided by longer, more nuanced inputs. For instance, ALLAM’s relatively high performance in Cultural Interpretation of proverbs highlights its capacity for deeper socio-cultural reasoning, likely attributed to its Arabic-centric training.
The results reveal a noticeable gap between multilingual and Arabic-native models. Models like Qwen and Mistral, which were trained on a wide range of languages, outperformed their Arabic-centric counterparts in overall accuracy. Their ability to generalize across dialectal variations, despite not being trained explicitly on Saudi dialects, reflects the benefits of exposure to diverse language data. These results align with observations from Palm [18] and related studies, where multilingual training correlated with higher robustness on non-standard Arabic input.
Nevertheless, Arabic-native models demonstrated strengths in specific areas. For example, ALLaM achieved high accuracy in tasks involving proverbs, while Jais performed relatively well in location-based recognition. These patterns suggest that domain-specific pretraining provides advantages in interpreting traditional or idiomatic expressions, even if it does not generalize as effectively across diverse question types. Such findings align with results from the AraDiCE benchmark [3], where models exhibited improved performance when test data resembled the training distribution.
To further demonstrate model differences, consider a Cultural Interpretation question based on the proverb:
{RLtext}
“اللي ما يعرف الصقر يشويه”
“He who doesn’t recognize the falcon cooks it”
While ALLAM accurately selected the metaphorical interpretation (“Those who lack knowledge may misuse valuable things”), other models like Jais and ACEGPT selected overly literal or generic options. This reflects ALLAM’s stronger grounding in socio-cultural expressions—a critical strength when applying LLMs in culturally sensitive settings.
In summary, our findings reveal a clear gap in current LLMs’ ability to comprehensively handle the diverse spectrum of Saudi Arabic dialects, especially in less standardized or culturally nuanced forms. While multilingual models demonstrate stronger generalization capabilities, Arabic-native models show targeted strengths in specific areas such as cultural interpretation. However, neither category achieves consistently strong performance across all dialects and task types. These insights underscore the need for more representative and balanced datasets, as well as tailored adaptation strategies, to improve model alignment with dialectal Arabic at the national level.
7 Ethical Considerations and Limitations
7.1 Ethical Considerations
This study utilizes publicly available data sources to evaluate the dialectal understanding of LLMs within the Saudi context. Care was taken to ensure that no private, identifiable, or sensitive information was included in the dataset. Given that language models may produce hallucinated or inaccurate outputs—particularly in low-resource or culturally nuanced scenarios—we stress that these outputs should not be used for critical decision-making without human oversight. The evaluation framework was carefully designed to minimize cultural bias and to avoid reinforcing stereotypes. Throughout the project, we maintained a commitment to transparency and responsible AI usage, acknowledging the inherent limitations of current LLMs when handling dialectal and culturally embedded language. Ultimately, our goal is to promote the development of fair, inclusive, and culturally aware language technologies.
7.2 Limitations
While the Absher benchmark represents a significant step toward evaluating LLMs on Saudi dialects and culture, the study is subject to several limitations. First, the evaluations were conducted under constrained computational resources, which limited our ability to explore multiple model configurations or conduct repeated trials. Second, all models were assessed using a fixed prompt structure (e.g., “Answer as if you are a Saudi…”), which may have influenced model behavior. Future work could explore the impact of prompt variation and instruction tuning. Third, although human evaluation was performed on a stratified sample to ensure dataset quality, a complete manual review of all 18,564 questions was not feasible due to the scale and diversity of the benchmark. Finally, while Absher covers various regional Saudi dialects, the representation of these dialects within the dataset is not perfectly balanced; for instance, Central and Western dialects are more prevalent. That highlights the need for more diverse and evenly distributed resources to comprehensively capture all Saudi dialectal variations in future research. Future research should leverage more extensive computational infrastructure, experiment with prompt engineering strategies, and incorporate broader expert reviews to further enhance the accuracy, reliability, and cultural validity of the benchmark evaluations.
8 Conclusion and Future Work
This paper introduced Absher, a novel benchmark designed to evaluate the capabilities of large language models (LLMs) in understanding the diverse dialects and rich cultural expressions across Saudi Arabia. Unlike existing Arabic benchmarks that center on Modern Standard Arabic (MSA) or treat dialects at a coarse level, Absher offers fine-grained, region-specific evaluation through over 18,000 carefully constructed multiple-choice questions. These questions span six task types—ranging from meaning identification to cultural interpretation—and are grounded in authentic dialectal content from various Saudi regions.
Our evaluation of several state-of-the-art LLMs revealed significant variation in performance across dialects, question types, and linguistic units. While some models performed well on tasks like location recognition and contextual usage, they struggled with cultural inference and binary reasoning, especially in low-context scenarios. Notably, multilingual models such as Qwen and Mistral outperformed Arabic-native models in general tasks, whereas the latter showed an advantage in interpreting culturally embedded expressions like proverbs.
Beyond technical performance, this work contributes to a broader sociolinguistic goal: fostering inclusive and culturally sensitive NLP systems. By systematically incorporating dialectal diversity and cultural context, Absher helps bridge existing evaluation gaps and supports the development of language technologies that reflect the lived linguistic realities of users.
Looking forward, future work will expand Absher to include additional dialects, question formats, and model types. We also plan to explore few-shot and instruction-tuned settings to assess model adaptability in low-resource scenarios. Integrating sociolinguistic features and region-specific datasets may further improve alignment with real-world language use. Through these efforts, we aim to promote more inclusive, robust, and context-aware Arabic NLP systems.
Data Availability
We provided several illustrative examples from our benchmark in this paper, and the complete dataset will be made publicly available for research purposes upon acceptance.
Declaration of competing interest
The authors declare that they have no competing interests.
Acknowledgment
Declaration of generative AI and AI-assisted technologies in the writing process: During the preparation of this work, we used GPT-4, an AI chatbot developed by OpenAI, to improve our written work. After using this tool/service, we reviewed and edited the content as needed and take full responsibility for the content of the publication.
Appendices
Appendix A Prompt Example
Appendix B Annotation Guideline
Figure 8: The interface designed using Google Sheets, which was shared with annotators to evaluate the questions.
Appendix C Detailed Performance
Model Content Type Question Type Accuracy Precision Recall F1-score
[Uncaptioned image] ALLaM Word Meaning 24.42% 28.08% 25.32% 10.48%
True/False 21.88% 20.45% 24.27% 9.21%
Fill-in-the-Blank 44.24% 39.54% 32.19% 28.19%
Contextual Usage 42.08% 48.53% 27.49% 19.46%
Cultural Interpretation 42.33% 38.99% 25.42% 15.86%
Location Recognition 67.44% 27.63% 25.04% 20.46%
Phrase Meaning 46.03% 48.09% 35.48% 24.49%
True/False 22.00% 5.55% 24.26% 9.03%
Fill-in-the-Blank 43.53% 32.71% 27.84% 23.35%
Contextual Usage 37.26% 50.58% 23.49% 17.40%
Cultural Interpretation 39.18% 9.82% 24.86% 14.07%
Location Recognition 65.50% 16.48% 24.92% 19.84%
Proverb Meaning 46.15% 35.00% 32.14% 25.40%
True/False 30.77% 29.17% 32.14% 29.09%
Fill-in-the-Blank 55.56% 38.51% 26.15% 21.53%
Contextual Usage 45.00% 47.22% 51.30% 42.04%
Cultural Interpretation 60.00% 58.79% 52.46% 52.99%
Location Recognition 53.25% 31.00% 30.36% 27.73%
[Uncaptioned image] LLaMA Word Meaning 25.98% 49.80% 25.49% 13.15%
True/False 27.24% 25.12% 24.93% 13.16%
Fill-in-the-Blank 42.64% 38.67% 27.45% 20.38%
Contextual Usage 43.55% 44.36% 22.98% 18.27%
Cultural Interpretation 41.99% 28.15% 20.12% 13.14%
Location Recognition 66.32% 30.85% 21.87% 20.66%
Phrase Meaning 42.86% 23.33% 24.88% 16.46%
True/False 29.56% 28.41% 26.60% 14.20%
Fill-in-the-Blank 45.91% 31.11% 27.16% 20.46%
Contextual Usage 38.33% 37.73% 23.43% 18.02%
Cultural Interpretation 40.69% 42.47% 26.00% 16.90%
Location Recognition 63.32% 24.41% 21.79% 20.35%
Proverb Meaning 30.77% 10.26% 33.33% 15.69%
True/False 32.50% 41.31% 38.33% 21.57%
Fill-in-the-Blank 56.79% 38.82% 28.12% 23.35%
Contextual Usage 40.00% 34.74% 26.56% 16.96%
Cultural Interpretation 46.25% 24.04% 25.81% 17.30%
Location Recognition 54.55% 13.64% 25.00% 17.65%
[Uncaptioned image] Jais Word Meaning 23.89% 30.87% 25.26% 10.02%
True/False 22.26% 22.82% 24.79% 9.26%
Fill-in-the-Blank 39.57% 35.41% 20.48% 12.29%
Contextual Usage 40.36% 33.62% 25.55% 15.98%
Cultural Interpretation 44.31% 52.65% 27.68% 20.96%
Location Recognition 67.57% 16.91% 24.97% 20.17%
Phrase Meaning 42.86% 14.29% 33.33% 20.00%
True/False 23.11% 40.92% 33.53% 12.75%
Fill-in-the-Blank 41.59% 51.99% 25.94% 16.68%
Contextual Usage 35.76% 31.09% 20.46% 11.77%
Cultural Interpretation 39.61% 45.27% 25.39% 16.10%
Location Recognition 65.72% 41.48% 25.15% 20.30%
Proverb Meaning 30.77% 10.26% 33.33% 15.69%
True/False 21.25% 7.08% 33.33% 11.68%
Fill-in-the-Blank 51.85% 12.96% 25.00% 17.07%
Contextual Usage 37.50% 9.38% 25.00% 13.64%
Cultural Interpretation 45.00% 11.25% 25.00% 15.52%
Location Recognition 54.55% 13.64% 25.00% 17.65%
[Uncaptioned image] Mistral Word Meaning 32.08% 29.91% 26.60% 22.09%
True/False 58.40% 58.15% 59.70% 56.56%
Fill-in-the-Blank 42.97% 10.74% 25.00% 15.03%
Contextual Usage 39.48% 9.87% 25.00% 14.15%
Cultural Interpretation 45.07% 11.27% 25.00% 15.53%
Location Recognition 69.77% 17.44% 25.00% 20.55%
Phrase Meaning 38.71% 27.44% 26.34% 24.64%
True/False 56.00% 58.28% 59.38% 55.36%
Fill-in-the-Blank 41.38% 35.31% 25.11% 14.82%
Contextual Usage 35.12% 7.04% 20.00% 10.41%
Cultural Interpretation 39.39% 9.85% 25.00% 14.13%
Location Recognition 58.70% 30.77% 46.51% 34.36%
Proverb Meaning 30.77% 11.11% 33.33% 16.67%
True/False 46.15% 30.00% 33.33% 31.58%
Fill-in-the-Blank 50.62% 12.97% 24.40% 16.94%
Contextual Usage 36.25% 9.18% 24.17% 13.30%
Cultural Interpretation 45.00% 11.25% 25.00% 15.52%
Location Recognition 53.85% 35.42% 35.00% 34.85%
[Uncaptioned image] Qwen Word Meaning 64.47% 35.52% 37.59% 35.81%
True/False 75.13% 31.78% 30.60% 30.82%
Fill-in-the-Blank 60.51% 47.27% 54.62% 48.08%
Contextual Usage 64.67% 52.41% 54.08% 52.96%
Cultural Interpretation 63.74% 40.79% 44.49% 41.28%
Location Recognition 51.45% 44.65% 56.56% 43.11%
Phrase Meaning 61.29% 36.27% 37.65% 36.12%
True/False 18.49% 32.23% 31.89% 11.45%
Fill-in-the-Blank 33.33% 25.04% 27.41% 16.58%
Contextual Usage 33.75% 24.97% 16.79% 13.41%
Cultural Interpretation 63.93% 39.62% 46.12% 40.16%
Location Recognition 51.97% 42.94% 51.74% 40.61%
Proverb Meaning 47.56% 24.38% 12.04% 16.12%
True/False 31.71% 23.13% 33.11% 13.64%
Fill-in-the-Blank 31.71% 20.00% 14.84% 15.69%
Contextual Usage 67.50% 59.91% 59.74% 57.64%
Cultural Interpretation 50.00% 24.40% 12.65% 16.67%
Location Recognition 31.71% 24.07% 8.02% 12.04%
[Uncaptioned image] ACEGPT Word Meaning 43.80% 26.37% 22.11% 21.07%
True/False 37.22% 19.86% 20.11% 14.96%
Fill-in-the-Blank 41.84% 35.63% 41.09% 33.47%
Contextual Usage 54.53% 47.50% 45.84% 44.14%
Cultural Interpretation 40.23% 33.23% 31.89% 26.85%
Location Recognition 40.28% 36.79% 38.93% 29.34%
Phrase Meaning 39.47% 23.94% 22.26% 21.26%
True/False 40.48% 27.03% 27.35% 20.72%
Fill-in-the-Blank 36.84% 34.81% 41.95% 29.54%
Contextual Usage 49.12% 48.07% 48.73% 41.73%
Cultural Interpretation 39.29% 35.71% 33.49% 26.63%
Location Recognition 36.95% 34.43% 41.52% 26.72%
Proverb Meaning 42.31% 44.70% 32.44% 34.03%
True/False 23.81% 41.67% 43.38% 23.64%
Fill-in-the-Blank 28.57% 26.67% 15.56% 19.64%
Contextual Usage 43.24% 50.75% 39.73% 32.26%
Cultural Interpretation 28.26% 49.38% 38.98% 29.33%
Location Recognition 38.67% 30.48% 37.80% 28.22%
Table 10: Evaluation Results of the Model’s Performance

### v2

@@
 Abstract
-As large language models (LLMs) become increasingly central to Arabic NLP applications, evaluating their understanding of regional dialects and cultural nuances is essential, particularly in linguistically diverse settings like Saudi Arabia.
-This paper introduces Absher, a comprehensive benchmark specifically designed to assess LLMs performance across major Saudi dialects.
-Absher comprises over 18,000 multiple-choice questions spanning six distinct categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
-These questions are derived from a curated dataset of dialectal words, phrases, and proverbs sourced from various regions of Saudi Arabia.
-We evaluate several state-of-the-art LLMs, including multilingual and Arabic-specific models.
-We also provide detailed insights into their capabilities and limitations.
-Our results reveal notable performance gaps, particularly in tasks requiring cultural inference or contextual understanding.
-Our findings highlight the urgent need for dialect-aware training and culturally aligned evaluation methodologies to improve LLMs performance in real-world Arabic applications.
+As large language models (LLMs) become increasingly central to Arabic NLP applications, their effectiveness in linguistically diverse settings, particularly regions with rich dialectal variation such as Saudi Arabia, remains underexplored.
+Existing evaluation paradigms tend to prioritize high-resource languages or Modern Standard Arabic (MSA), overlooking regional linguistic and cultural specificities.
+This leads to performance limitations and cultural biases in real-world deployments.
+To address this gap, we introduce Absher, the first comprehensive and fine-grained benchmark designed to assess the understanding of LLMs regarding Saudi dialects and their embedded cultural nuances.
+Absher consists of over 18,000 multiple choice questions derived from a curated dataset of dialectal words, phrases, and proverbs sourced from five major Saudi regions.
+The benchmark spans six task categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition, enabling multifaceted evaluation across both linguistic and cultural dimensions.We perform zero-shot evaluations on six state-of-the-art open LLMs: ALLaM, LLaMA, Jais, Mistral, Qwen, and AceGPT.
+Our results reveal substantial performance variability across dialects and question types.
+Qwen achieved the highest overall accuracy, excelling in word-level questions (63%), while ALLaM outperformed others in the interpretation of proverbs (48% accuracy).
+All models struggled with content from underrepresented dialects, particularly Southern and Eastern variants, and with context-free True/False questions, highlighting weaknesses in dialect grounding and binary reasoning.
+These findings demonstrate the need for dialect-aware training and culturally aligned evaluation.
+We position Absher as a critical step towards more equitable and effective LLM development for real-world Arabic applications.
 keywords:
-Large Language Models, Natural Language Processing , Dialects Understanding
+Large Language Models, Natural Language Processing , Cultural aware LLMs, Dialects Understanding
 1 Introduction
-Recent breakthroughs in the area of language processing have shown that contemporary computational models can perform a variety of tasks, including text production, question answering, and translation between languages [1, 2].
-Most of the models, though, have been developed with a prevailing focus on English and Modern Standard Arabic (MSA), limiting their performance in domains with linguistic and cultural diversity, such as Saudi Arabia.
-These systems tend to find it difficult to interpret expressions related to local identity like idiomatic expressions, cultural practices, and dialectal tones that are underrepresented in the data used to train them [3, 4].
-Saudi Arabia, as depicted in Figure 1, have a unique and rich linguistic landscape characterized by a significant variety of regional dialects.
-These include Central, Western, Southern, Eastern, and Northern, each with distinct features in pronunciation, vocabulary, and grammar [5].
-Beyond phonetic and lexical differences, these dialects are intimately linked to region-specific proverbs, traditional expressions, and patterns of social interaction, forming a fundamental part of ordinary conversation.
-While essential, these features are rarely incorporated into existing language resources or evaluation tools [6, 7].
-Therefore, there is growing academic interest in exploring ways to enable the proper interpretation of such dialectal and cultural features in computational models.
-Figure 1: A map of Saudi Arabia showing the different regions and their dialectical variations.
-Beyond improving technical accuracy, interpreting Saudi dialects also has broader sociolinguistic significance.
+Recent breakthroughs in the area of language processing have shown that contemporary computational models can perform a variety of tasks, including text production, question answering, and machine translation [Brown2020, OpenAI2023].
+These powerful computational models are becoming indispensable tools across various sectors, from education and healthcare to public services and entertainment.
+However, a critical challenge remains: the vast majority of these models, along with their foundational training data and evaluation benchmarks, are developed with a prevailing focus on high-resource languages like English or, within the Arabic-speaking world, Modern Standard Arabic (MSA).
+This limits their performance and applicability in domains characterized by significant linguistic and cultural diversity, particularly those with rich regional dialects [Baly2023, alwajih2025palm].
+Language is inherently cultural; its full comprehension demands reasoning beyond literal meaning to grasp social implications and regional specificities.
+Saudi Arabia exemplifies this challenge with its rich landscape of regional dialects.
+These dialects are not merely linguistic variants, but embody unique cultural practices, idiomatic expressions, and proverbs critical to authentic communication [al2018suar].
+Despite this profound linguistic and cultural diversity, existing Arabic natural language processing (NLP) benchmarks and large language models (LLMs) training data largely overlook these fine-grained Saudi dialectal and cultural nuances [ayash2025saudiculture, Baly2023].
+Therefore, models struggle to interpret local identity, leading to significant performance gaps and potentially biases against underrepresented languages.
+To address this critical gap and rigorously assess LLMs’ cultural and linguistic competence in this domain, we introduce Absher, a novel, comprehensive, and first-of-its-kind benchmark.
+Absher offers an evaluation framework built on a curated dataset of authentic Saudi dialectal content, sourced from diverse regions across the Kingdom.
+It comprises $18,564$ multiple-choice questions, systematically generated to cover six distinct categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
+This multi-faceted design facilitates a deep, granular assessment of LLMs’ capabilities that moves beyond superficial linguistic understanding toward genuine cultural competence.
+Beyond technical accuracy, the ability of LLMs to process and respond to Saudi dialects carries broader sociolinguistic significance.
 These dialects are not only linguistic systems but also carriers of regional identity, social norms, and cultural heritage.
-Ensuring that LLMs can process and respond to them fairly contributes to linguistic inclusivity and mitigates the marginalization of underrepresented communities.
+Ensuring their equitable representation in AI systems is crucial for linguistic inclusion and mitigating the marginalization of underrepresented communities.
 Such efforts are vital for the development of equitable AI systems and for informing culturally aware language technologies and public policy.
-Current benchmarks for Arabic language models are based primarily on MSA or generalized applications [8, 9], with very little attention paid to dialect and cultural representation.
-This under-representation affects the fairness as well as the effectiveness of LLMs when utilized by Arabic speakers in real-world situations.
-In particular, Saudi dialects and cultural elements are less represented in training sets and test sets, leading to inferior performance on culturally contextualized tasks like the interpretation of idiomatic expressions or answering appropriately in contextually sensitive contexts [10, 11].
-To address this critical gap, our work introduces Absher, a novel benchmark specifically designed to assess the performance of large language models (LLMs) on the unique and fine-grained linguistic and cultural nuances of Saudi dialects.
-As illustrated in Figure 2, Absher provides a comprehensive evaluation framework built upon a curated set of dialectal contents—including words, phrases, and proverbs—representing diverse regions across Saudi Arabia.
-The benchmark comprises a total of 18,564 multiple-choice questions, systematically generated to cover six distinct types: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
-Figure 2: Overview of our benchmark.
-In our work, we aim to provide answers to these key questions.
-First, we investigate the ability of LLMs to interpret Saudi dialects and cultural expressions by creating a new benchmark dataset that captures the nuances of both.
-Second, we identify the main challenges LLMs face in this context through a detailed error analysis of their performance on this benchmark.
-Finally, we present a comparative analysis of different LLMs, evaluating their strengths and weaknesses in handling Saudi dialect and culture-specific tasks.
-Our work addresses the aforementioned research questions through the following core objectives:
-  • Benchmark Development: Construct a culturally grounded benchmark that systematically evaluates LLMs across diverse Saudi dialects.
-The benchmark includes multiple question formats targeting both linguistic comprehension and cultural awareness, enabling fine-grained assessment.
-  • Performance Evaluation: Conduct a rigorous evaluation of several LLMs using the constructed benchmark to assess their capabilities in interpreting Saudi dialects.
-The analysis identifies their strengths, limitations, and areas that require further research and refinement.
+Figure 1: Overview of Absher, covering all regions of Saudi Arabia (five specific regions and a general category) and presents representative examples from its three types of content: words, phrases, and proverbs.
+The benchmark also spans all six question types: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
+In this work, our objective is to provide answers to key research questions concerning LLMs’ capabilities in interpreting Saudi dialects, the specific challenges they face, and a comparative analysis of their performance.
+Our goal is twofold: (1) to develop a novel, fine-grained benchmark that systematically evaluates LLMs across Saudi dialects and linguistic units; and (2) to conduct a comprehensive evaluation of state-of-the-art LLMs, analyzing their performance, weaknesses, and areas for improvement.
+Our key contributions are as follows:
+  • Benchmark Development: We introduce Absher, the first large-scale benchmark tailored to Saudi dialects, featuring over 18,000 multiple-choice questions.
+These questions span six task categories and are grounded in authentic dialectal words, phrases, and proverbs from five major Saudi regions, addressing a critical gap in Arabic NLP evaluation.
+Figure  1 illustrates the composition of Absher across different regions and categories.
+  • Methodological Framework: We propose a scalable pipeline that combines prompt engineering with automated question generation using LLMs, followed by a rigorous multi-stage human validation process.
+This ensures high linguistic and cultural fidelity, making Absher a reliable tool for dialectal evaluation.
+  • Model Evaluation and Analysis: We conduct a comprehensive evaluation of multilingual and Arabic-specific LLMs on the Absher benchmark.
+Our analysis reveals significant performance variability across dialects, task categories, and linguistic units.
+We also investigate the impact of model size, language orientation, and architecture, offering insights into the limitations and strengths of current LLMs in dialect-sensitive settings.
+This paper is structured as follows: Section 2 reviews the existing literature on Arabic dialects in LLMs.
+Section 3 details the methodology for constructing the Absher benchmark, including data collection, pre-processing, prompt design, question generation, and human evaluation.
+Section 4 presents the experimental setup and results of our LLM evaluations.
+Section 5 provides a detailed analysis of our findings, including comparative performance across models, content types, and question types.
+Finally, Section 6 discusses the broader implications of our results and outlines promising directions for future research.
 2 Related Work
 2.1 Arabic Dialects in LLMs
-Recent advances in Arabic NLP have led to increased attention toward evaluating large language models (LLMs) on dialectal and cultural understanding.
-Several benchmarks have been proposed to assess model performance in these dimensions.
-For instance, AraDiCE [7] introduced a benchmark targeting dialectal and cultural understanding across Levantine, Egyptian, and Gulf Arabic, revealing critical limitations in generation, translation, and reasoning capabilities.
-Calibration-Arabic-PLMs [12] focused on model confidence calibration in dialect identification tasks, highlighting inconsistencies between model certainty and actual performance.
-Health-LLM-Arabic [13] evaluated LLMs like GPT-4 on health-related claims involving Saudi, Egyptian, and Moroccan dialects, and proposed the Cultural Sensitivity Score to quantify dialect adaptation in clinical scenarios.
-Complementing these evaluation-focused efforts, a number of datasets and models have been introduced to enhance dialect and culture awareness in Arabic LLMs.
-Dallah [14] presented a dialect-aware multimodal assistant fine-tuned across six Arabic dialects.
-AraReasoner [15] provided a benchmarking suite for assessing the reasoning capabilities of LLMs across a wide range of Arabic NLP tasks.
-Jawaher [16] curated a multidialectal dataset of Arabic proverbs designed to test LLMs’ understanding of figurative and culturally grounded expressions.
-Similarly, CARE [17] introduced a multilingual benchmark with human preference ratings aimed at evaluating cultural alignment in Arabic and Chinese contexts.
-Palm [18] contributed a linguistically and culturally diverse dataset spanning 22 Arab countries, with instructions formulated in both MSA and regional dialects to support comprehensive evaluation.
-In the multimodal space, several resources have emerged to test vision-language capabilities in culturally rich Arabic settings.
-Peacock [19] introduced a family of Arabic multimodal LLMs and released Henna, a benchmark designed to assess cultural competence.
-JEEM [20] focused on visual question answering and image captioning across four Arabic dialects, using culturally specific imagery.
-Lastly, Swan and ArabicMTEB [21] offered a family of Arabic-centric embedding models and a robust benchmark suite for evaluating cross-lingual, multi-dialectal, and multi-domain embedding quality.
+Recent efforts in Arabic NLP have increasingly emphasized dialectal and cultural aspects in evaluating LLMs.
+Several benchmarks have emerged to measure model performance in these dimensions.
+For example, AraDiCE [Baly2023] assessed LLMs on dialectal and cultural understanding across Levantine, Egyptian, and Gulf Arabic, highlighting limitations in generation and reasoning.
+Calibration-Arabic-PLMs [Al-Laith2025] examined model confidence in dialect identification, revealing misalignment between confidence and accuracy.
+Health-LLM-Arabic [alharbi-etal-2025-evaluating] introduced a Cultural Sensitivity Score to evaluate model responses to health-related claims in Saudi, Egyptian, and Moroccan dialects.
+In addition, to these evaluation efforts, several datasets and models have been designed to enhance dialect awareness.
+Dallah [alwajih2024dallah] developed a multimodal assistant fine-tuned on six dialects, while AraReasoner [hasanaath2025arareasoner] proposed a benchmark suite for evaluating Arabic reasoning capabilities.
+Jawaher [magdy2025jawaher] focused on figurative language through a collection of Arabic proverbs, and CARE [guo2025care] evaluated cultural alignment via human preferences in Arabic and Chinese contexts.
+Palm [alwajih2025palm] contributed a linguistically diverse dataset covering 22 Arab countries, combining MSA and dialectal instructions.
+In the multimodal space, Peacock [alwajih2024peacock] released the Henna benchmark to test cultural competence, and JEEM [kadaoui2025jeem] introduced dialect-aware visual tasks involving culturally grounded images.
+ArabicMTEB and Swan [bhatia2024swan] offered resources for evaluating cross-lingual, dialectal, and domain-specific Arabic embeddings.
 2.2 Saudi Dialects in LLMs
-Despite growing interest in Arabic NLP, few studies have explicitly focused on evaluating large language models (LLMs) within the context of Saudi dialects.
-One of the first targeted efforts is SaudiCulture [6], a benchmark designed to assess LLMs’ cultural competence across five major regions of Saudi Arabia.
-Evaluations of models such as GPT-4, LLaMA 3.3, FANAR, Jais, and AceGPT revealed notable performance degradation when addressing region-specific or culturally grounded questions.
-Similarly, AL-QASIDA [22] evaluated nine LLMs across eight Arabic dialects, including Saudi Arabic, analyzing generation quality through fidelity, fluency, and diglossia metrics.
-The study highlighted the impact of pre-training data and suggested methods for dialect-aware model adaptation.
-In broader benchmarks such as Dallah [14] and Palm [18], Saudi dialects were included alongside other regional variants, though without a focused or fine-grained analysis of Saudi-specific linguistic features.
-Beyond direct evaluation, other efforts have explored data creation and translation to support dialectal Arabic processing.
-For instance, LLM-MT-Data [23] employed LLMs to translate dialectal Arabic—sourced from the Saudi Audio Dataset for Arabic (SADA)—into Modern Standard Arabic, demonstrating the potential of LLMs in generating high-quality parallel data.
-Similarly, AraBench [24] tackled dialect-specific machine translation using fine-tuning and back-translation techniques to improve Arabic–English translation quality.
-While these initiatives mark important progress, most prior work has either concentrated on broader Arabic dialects or MSA, without a dedicated focus on the linguistic and cultural intricacies of Saudi dialects.
-Dialects such as Central, Western, and Southern are deeply intertwined with region-specific expressions, proverbs, and cultural practices that remain underexplored in current benchmarks.
-Notably, no prior study has systematically examined Saudi dialects across diverse linguistic units—such as words, phrases, and proverbs—or assessed LLMs’ ability to interpret them in context.
-This gap underscores the need for a benchmark like Absher, which provides a structured, culturally grounded evaluation of LLMs within the Saudi dialect landscape.
+Despite growing interest in Arabic dialects, Saudi-specific evaluations remain limited.
+SaudiCulture [ayash2025saudiculture] evaluated five LLMs across five Saudi regions, revealing performance drops on culturally grounded content.
+AL-QASIDA [robinson2024qasida] examined nine LLMs across eight dialects—including Saudi Arabic—using fidelity and diglossia metrics, and discussed pretraining implications.
+Other benchmarks, such as Dallah [alwajih2024dallah] and Palm [alwajih2025palm], included Saudi dialects but lacked focused analysis.
+LLM-MT-Data [Elmogtaba2024] used LLMs to translate Saudi speech data into MSA, and AraBench [sajjad2020arabench] explored dialect-to-English translation with fine-tuning and back-translation.
+A notable data-centric effort is the KIND dataset [yamani2024kind], which collected fine-grained dialectal Arabic through a community-driven competition.
+KIND includes parallel corpora and Q&A spanning 29 Saudi dialects, with city-level granularity.
+However, it is primarily designed for data collection and does not offer a structured evaluation of LLMs’ contextual understanding, a gap Absher seeks to fill.
+A key point of contrast is with SaudiCulture [ayash2025saudiculture], which introduced a pioneering cultural benchmark.
+However, it is limited to 441 questions focused mainly on general knowledge and cultural themes predominantly in MSA.
+In contrast, Absher provides a broader and more targeted evaluation with a significantly larger volume of questions.
+It is uniquely designed for a fine-grained, dialect-sensitive assessment across Saudi dialects, focusing on the nuanced understanding of words, phrases, and proverbs.
+Unlike SaudiCulture, which emphasizes high-level cultural reasoning, Absher enables a deep, multi-category evaluation drawn from region-specific vernacular data, making it a more comprehensive and regionally grounded resource for advancing Arabic NLP.
+While these studies offer valuable insights, none provide a comprehensive benchmark that systematically evaluates LLMs on Saudi dialects across linguistic units such as words, phrases, and proverbs.
+Existing work often overlooks the sociocultural specificity of regional expressions.
+In contrast, Absher introduces a fine-grained, culturally contextualized benchmark that spans multiple Saudi dialects and domains, enabling detailed evaluation of LLMs’ dialectal and cultural understanding.
+It further differentiates itself through real-world question design, multi-unit coverage, and region-aware analysis of model behavior.
 3 Constructing Absher Benchmark
 To assess the ability of computational systems to understand the dialectal and cultural nuances specific to Saudi Arabia, we introduce the Absher Benchmark.
-This evaluation framework is built upon a curated set of dialectal contents—including words, phrases, and proverbs—representing diverse regions across Saudi Arabia.
-The development process followed a structured pipeline encompassing data collection, task formulation, prompt engineering, automated question generation, and expert-based review as seen in Figure  3.
+This evaluation framework is built upon a curated set of dialectal contents, including words, phrases, and proverbs, representing diverse regions across Saudi Arabia.
+The development process followed a structured pipeline encompassing data collection, pre-processing, prompt engineering, automated question generation, and expert-based review as seen in Figure 2.
 Figure 2: The overall pipeline of constructing Absher benchmark.
@@
 To evaluate the ability of computational models to interpret Saudi dialects and culturally embedded expressions, we introduce Absher, a benchmark grounded in authentic regional content.
-It is built from a curated collection of dialectal items—including words, phrases, and proverbs—representing linguistic diversity across Saudi Arabia.
-The benchmark construction followed a structured pipeline encompassing data collection, content classification, task formulation, prompt engineering, automated question generation, and expert validation, as illustrated in Figure 3.
+It is built from a curated collection of dialectal items, including words, phrases, and proverbs, representing linguistic diversity across Saudi Arabia.
 The core of the dataset was collected from the Moajam website [^1] https://ar.mo3jam.com/, a digital repository that aggregates dialectal vocabulary across Arabic regions.
 We scraped and organized entries into three categories: words, phrases, and proverbs, each annotated with its meaning, dialect label, contextual usage, and translation.
-This process ensured comprehensive coverage of both literal and figurative expressions relevant to regional Saudi Arabic.
-Words refer to vocabulary items specific to Saudi dialects, often including slang or regional variants not found in Modern Standard Arabic—for example, ”Rabshah”.
-Phrases encompass idiomatic or conversational expressions commonly used across regions, such as ”Allah Yta’anni”.
-Proverbs are culturally rooted sayings that convey implicit meanings, requiring contextual and cultural understanding—e.g., ”Ally ma ya’arf alsger yshwaeh”.
-The raw collection initially consisted of 5,483 items, which were refined through preprocessing and filtering, resulting in 3,094 unique contents used to generate over 18,000 multiple-choice questions across six task types.
+This process ensured a comprehensive coverage of both literal and figurative expressions relevant to regional Saudi Arabic.
+  • Words refer to vocabulary items specific to Saudi dialects, often including slang or regional variants not found in Modern Standard Arabic.
+  • Phrases encompass idiomatic or conversational expressions commonly used across regions.
+  • Proverbs are culturally rooted sayings that convey implicit meanings, requiring contextual and cultural understanding.
+The raw collection initially consisted of 5,483 items, which were refined through pre-processing and filtering, resulting in 3,094 unique contents used to generate over 18,000 multiple-choice questions across six task types.
 Table 1 summarizes the dataset statistics.
+Table 1: Content statistics before and after pre-processing and the number of generated questions by category.
 Category # Contents (Raw) # Contents (Filtered) # Generated Qs
@@
 Total 5,483 3,094 18,564
-Table 1: Content statistics before and after preprocessing and the number of generated questions by category.
 3.2 Pre-processing Steps
@@
 To rigorously assess the capabilities of LLMs in interpreting Saudi dialects, we employed a structured prompt design strategy tailored to dialectal understanding.
-Each prompt template included four core components: a task description, a dialect-based question, multiple-choice answer options, and a designated correct answer (see Table 2).
-Prompts were derived from dialectal contents—words, phrases, and proverbs—each annotated with its dialect label, meaning, and a contextual usage example.
+Each prompt template included four core components: a task description, a dialect-based question, multiple-choice answer options, and a designated correct answer.
+Table 2 provides an illustration.
+The prompts were derived from dialectal contents, such as words, phrases, and proverbs.
+Each annotated with its dialect label, meaning, and a contextual usage example.
 The initial prompt generation followed a simple formulation process, but several issues emerged.
-Correct answers frequently appeared in fixed positions (e.g., always option A), True/False questions were heavily biased toward “True,” and some questions included incorrect dialectal labels.
+Correct answers frequently appeared in fixed positions (e.g., always option A), True/False questions were heavily biased toward “True”, and some questions included incorrect dialectal labels.
 These inconsistencies introduced unintended biases, undermining the fairness and reliability of the benchmark.
 To address these limitations, we refined the prompt design by introducing structural controls.
-Correct answer positions were randomized to prevent positional bias, and the distribution of True/False items was balanced to ensure fairness.
+The correct answer positions were randomized to prevent positional bias, and the distribution of True/False items was balanced to ensure fairness.
 Dialect annotations were reviewed and validated by linguistic experts to confirm their accuracy.
+Crucially, we implemented a rigorous two-step filtering strategy to handle abnormal model outputs from GPT-4o:
+  • Automated Filtering (Formatting and Structure): All generated outputs were subjected to a validation script to check for strict structural integrity.
+This script automatically flagged and discarded any questions that contained malformed JSON output, failed to include exactly four distinct multiple-choice options, or did not designate a clear correct answer.
+This ensured only well-structured, testable items proceeded to the next phase.
+  • Manual Curation (Hallucination and Semantic Errors): A team of two native Saudi Arabic speakers and linguistic experts manually reviewed a sample of automatically validated questions.
+Their primary task was to filter out semantic errors and hallucinations, specifically removing questions where the generated options were culturally inaccurate, the question’s premise contradicted the source material (meaning or context), or the designated correct answer was objectively wrong.
 These refinements significantly improved the robustness, validity, and representativeness of the final set of multiple-choice questions, establishing Absher as a dependable tool for dialect-sensitive LLMs evaluation.
 For the complete prompt template, please refer to Appendix A.
+Table 2: Template for question formulation
 Component Description
@@
 Correct Answer: The correct answer to the question.
-Table 2: Template for Question Formulation
 3.4 Question Generation
@@
 Compared to smaller or monolingual models, GPT-4o also exhibited greater consistency and fewer hallucinations—qualities essential for ensuring the validity of benchmark items.
-Each cleaned and preprocessed content item—whether a word, phrase, or proverb—was programmatically paired with its dialect label, meaning, and usage example to construct a complete input prompt.
+Each cleaned and pre-processed content item, whether a word, phrase, or proverb, was programmatically paired with its dialect label, meaning, and usage example to construct a complete input prompt.
 This setup enabled GPT-4o to generate questions that evaluated not only surface-level understanding but also deeper comprehension of dialectal and cultural nuances.
@@
   • Location Recognition: Identifying the geographic or regional origin associated with the dialectal term.
-This systematic generation approach ensured comprehensive coverage of the linguistic and cultural features embedded in Saudi dialects.
-By applying the process across all content categories—words, phrases, and proverbs—a total of 18,564 multiple-choice questions were generated.
-These questions form the core of the Absher benchmark and enable a rigorous, multifaceted evaluation of LLM performance.
+This systematic generation approach ensured comprehensive coverage of the linguistic and cultural characteristics embedded in the Saudi dialects.
+By applying the process across all content categories, words, phrases, and proverbs, a total of 18,564 multiple-choice questions were generated.
+These questions form the core of the Absher benchmark and allow for a rigorous, multifaceted evaluation of LLM performance.
 Table 3 shows the distribution of questions across regional dialects, highlighting the benchmark’s broad linguistic representation.
+The distribution of questions across regional dialects reflects the linguistic landscape of Saudi Arabia.
+The Najdi dialect, spoken predominantly in the Central region, and the Hejazi dialect, prevalent in the Western region, are spoken by a larger portion of the population and receive extensive exposure through media and online platforms.
+This widespread usage provides a richer pool of linguistic material.
+By contrast, the Southern, Northern, and Eastern dialects have a more localized presence, with fewer readily accessible resources.
+Despite these differences, the benchmark includes all major dialects, thus capturing the full spectrum of Saudi Arabic and supporting a comprehensive evaluation of LLM performance.
+Table 3: Number of questions per regional dialects in the Absher benchmark.
 Dialect # Qs
-Center 5,652
+Central 5,652
 Western 5,262
@@
 General Saudi Terms 4,338
-Table 3: Number of questions per regional dialects in the Absher benchmark.
-Figure 4 illustrates the six question types included in the benchmark.
+Figure 3 illustrates the six types of questions included in the benchmark.
 Together, these types enable a well-rounded evaluation of model capabilities, covering meaning interpretation, cultural and regional understanding, and recognition of local expressions.
@@
 3.5 Human Evaluation
-To validate the reliability, correctness, and cultural appropriateness of the questions in the Absher benchmark, we conducted a structured human evaluation.
+To validate the reliability, correctness, and cultural appropriateness of the questions in the Absher benchmark, we performed a structured human evaluation.
 This step was essential to ensure that the automatically generated content adhered to linguistic standards and dialectal relevance, especially given the complexity and regional variability of Saudi dialects.
@@
 To test attentiveness and accuracy, the pilot set included quality control questions with plausible yet incorrect answers.
-Annotators who misclassified these were given follow-up clarifications to align their understanding with the evaluation standards.
-This step significantly improved consistency and judgment reliability in the formal evaluation.
-The main evaluation involved a stratified random sample comprising 10% of the benchmark contents—Words, Phrases, and Proverbs—across all six question types.
+Annotators, who misclassified these, were received follow-up clarifications to align their understanding with the evaluation standards.
+This step improved consistency and reliability of the judgment in the formal evaluation.
+The main evaluation involved a stratified random sample comprising 10% of the benchmark contents, i.e., Words, Phrases, and Proverbs, across all six question types.
 This resulted in 600 annotated questions.
@@
 Additionally, we computed the overall alignment between model-generated answers and human judgments, which reached 93.03%, confirming the reliability and consistency of the benchmark’s human evaluation process.
-Evaluation Meaning True/False FillB Cont.Usage Cult.Interpretation Loca. Recognition
-Q-Quality 0.97 0.95 0.71 0.72 0.83 0.85
-A-Correctness 0.97 0.87 0.91 0.88 0.94 0.92
-Table 4: Cohen’s Kappa scores across the two evaluated criteria. FillB, Cont., Cult., and Loca. refer to “contextual”, “cultural”, “location”, and “Fill-in-the-Blank”.
+Table 4: Cohen’s Kappa scores for question quality and answer correctness across different question types. Bold denotes the top scores.
+Q-Type Q-Quality A-Correctness
+Meaning 0.97 0.97
+True/False 0.95 0.87
+Fill-in-the-Blank 0.71 0.91
+Contextual Usage 0.72 0.88
+Cultural Interpretation 0.83 0.94
+Location Recognition 0.85 0.92
 4 Experiments
-We evaluated a set of open-source LLMs: LLaMA-3 8B Instruct [25], Jais-13B [26], ALLaM-7B [27], Mistral-7B [28], Qwen2.5-7B Instruct [29], and AceGPT-7B-chat [30].
+We evaluated a set of open-source LLMs: LLaMA-3 8B Instruct [Touvron2024], Jais-13B [sengupta2023], ALLaM-7B [bari2025allam], Mistral-7B [jiang2023], Qwen2.5-7B Instruct [yang2024qwen2p5], and AceGPT-7B-chat [huang2023acegpt].
 These models were chosen based on three main criteria: open availability, Arabic language support, and competitive performance across a range of NLP tasks.
@@
 Since the benchmark is composed exclusively of multiple-choice questions, we employed four standard classification metrics: Accuracy, Precision, Recall, and F1 Score.
-These metrics offer a comprehensive and balanced evaluation of model performance, in line with prior work on zero-shot LLMs assessment for low-resource languages [31].
+These metrics offer a comprehensive and balanced evaluation of model performance, in line with prior work on zero-shot LLMs assessment for low-resource languages [lowres_llm_eval_2025].
+Table 5: Average evaluation results of models performance.
+Bold denotes the best result achieved across all models for each content type, while underlining indicates the best-performing content type for each individual model.
 Model Content Type Accuracy Precision Recall F1-score
@@
 Proverb 34.81% 37.93% 33.15% 27.51%
-Table 5: Average Evaluation Results of Models Performance
 4.1 Results
-Table 5 presents the performance of the six evaluated models—ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT—across different content types, namely words, phrases, and proverbs.
-The results reveal noticeable variations in performance depending on both the linguistic content and the complexity of the task [^2] The full set of results is provided in Appendix C..
+Table 5 presents the performance of the six evaluated models, such as ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT, across different content types, namely words, phrases, and proverbs.
+The results reveal noticeable variations in performance depending on both the linguistic content and the complexity of the task [^2] The full set of results is provided in Appendix D..
 For word-level content, the Qwen model demonstrated the strongest capabilities, achieving an accuracy of 63.00% and an F1-score of 41.18%.
 These results indicate Qwen’s effectiveness in handling isolated lexical items, suggesting its advantage in processing tasks that rely on clear, context-independent meanings.
+To contextualize the reported performance, it is important to note the chance-level baseline for this benchmark.
+Since most multiple-choice questions contain four options, random guessing corresponds to 25% accuracy (and 50% for True/False items).
+Therefore, the observed performance range of approximately 40–60% accuracy across models is substantially above chance, highlighting both the non-trivial capability of the evaluated LLMs and the inherent difficulty of the dialect- and culture-sensitive questions in Absher.
 In contrast, phrase-level tasks presented a more competitive landscape.
@@
 The set of models includes Arabic-specific and multilingual LLMs, spanning different sizes and scales, from 7B to 13B parameters.
-Model Size Language Accuracy [Uncaptioned image] Qwen2.5 Instruct 1st 7B Multilingual 50.35% [Uncaptioned image] Mistral 2nd 7B Multilingual 45.93% [Uncaptioned image] ALLaM 3rd 7B Arabic 43.78% [Uncaptioned image] LLaMA-3 Instruct 8B Multilingual 43.34% [Uncaptioned image] Jais 13B Arabic 42.14% [Uncaptioned image] AceGPT-chat 7B Arabic 38.88%
-Table 6: Models Overview, and Average accuracy
+Table 6: Overview of the evaluated models, highlighting their size, supported languages, and average accuracy achieved in the experiments.
+The best results are listed from highest to lowest, with the three best results in bold.
+Model Size Language Accuracy
+[Uncaptioned image] Qwen2.5 Instruct 7B Multilingual 50.35%
+[Uncaptioned image] Mistral 7B Multilingual 45.93%
+[Uncaptioned image] ALLaM 3rd 7B Arabic 43.78%
+[Uncaptioned image] LLaMA-3 Instruct 8B Multilingual 43.34%
+[Uncaptioned image] Jais 13B Arabic 42.14%
+[Uncaptioned image] AceGPT-chat 7B Arabic 38.88%
 While it is often assumed that larger parameter counts lead to better performance, our results demonstrate that accuracy does not scale linearly with size.
@@
 5.2 Examination of Model Responses by Region
-Our evaluation of Large Language Models (LLMs) responses to region-specific linguistic questions, detailed in Figure 5, reveals considerable performance variability influenced by both the geographic origin of expressions and the nature of the linguistic task.
+Our evaluation of LLMs responses to region-specific linguistic questions, detailed in Figure 4, reveals considerable performance variability influenced by both the geographic origin of expressions and the nature of the linguistic task.
 Notably, we observed strong performance in the General category for a widely known proverb, with five out of six models providing correct answers.
@@
 Figure 6: Average Accuracy for Different Models by Question Type
-Content Type: Proverb
-Question Type: Meaning
-Region: General
-Question Example {RLtext} المهمة: اختر إجابة واحدة فقط دون أي ملاحظات إضافية.
-السؤال: ماذا يعني المثل إذا حجت البقرة على قرونها؟
-الخيارات:
-أ) شيء يمكن تحقيقه إذا تمت الظروف بشكل جيد
-ب) شيء مستحيل أو غير قابل للتحقيق
-ج) شيء يحدث في حالة نادرة
-د) شيء يجب أن يتم في الوقت المحدد
-الإجابة الصحيحة: ب
-Models’ Answer: [Uncaptioned image] ALLaM: B , [Uncaptioned image] LLaMA: B, [Uncaptioned image] Jais: B, [Uncaptioned image] Mistral: C, [Uncaptioned image] Qwen: B, [Uncaptioned image] ACEGPT: B
-Table 7: Model answers to a proverb meaning question in the general dialect.
-Content Type: Proverb
-Question Type: Fill-in-the-Blank
-Region: Southern
-Question Example {RLtext} المهمة: اختر إجابة واحدة فقط دون أي ملاحظات إضافية.
-السؤال: خطط في _ _ _ واقبص في حيد.
-الخيارات:
-أ) التراب
-ب) الورق
-ج) الماء
-د) الصخر
-الإجابة الصحيحة: ج
-Models’ Answer: [Uncaptioned image] ALLaM: D , [Uncaptioned image] LLaMA: B, [Uncaptioned image] Jais: D, [Uncaptioned image] Mistral: D, [Uncaptioned image] Qwen: A, [Uncaptioned image] ACEGPT: D
-Table 8: Model answers to a fill-in-the-blank proverb in the southern dialect.
-Content Type: Word
-Question Type: Location Recognition
-Region: Central
-Question Example {RLtext} المهمة: اختر إجابة واحدة فقط دون أي ملاحظات إضافية.
-السؤال: أين يُستخدم المصطلح أويلاوه بشكل أكثر شيوعاً؟
-الخيارات:
-أ) القصيم
-ب) الرياض
-ج) جدة
-د) الدمام
-الإجابة الصحيحة: أ
-Models’ Answer: [Uncaptioned image] ALLaM : A , [Uncaptioned image] LLaMA : B, [Uncaptioned image] Jais : A, [Uncaptioned image] Mistral : B, [Uncaptioned image] Qwen : C, [Uncaptioned image] ACEGPT : D
-Table 9: Model answers to a Location Recognition Word question in the Central dialect.
 5.4 Case Study of LLMs Responses Across Question Types and Dialects
@@
 The results indicate that regionally common and widely used expressions are more accurately interpreted by LLMs.
-For example, in the general category, the proverb:
-{RLtext}
-“إذا حجت البقرة على قرونها”
-“If the cow performs Hajj on its horns” (Table 7) was correctly understood by five out of six models: ALLaM, LLaMA, Jais, Qwen, and ACEGPT all provided the correct answer, while only Mistral failed.
-This proverb is relatively widespread, and its meaning—“something impossible to achieve”—is often documented in public sources, likely increasing its representation in model training data.
-In contrast, proverbs that are more regionally isolated or culturally specific, such as the Southern fill-in-the-blank proverb :
-{RLtext}
-“خطط في ماء واقبص في حيد”
-“Plan in water and pinch in a rock” (Table 8), saw a complete failure across all six models.
-None of the models—ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT selected the correct option, suggesting a consistent gap in exposure to underrepresented regional idioms, especially those originating from areas with fewer accessible textual resources.
+For example, in the general category, the proverb: ”Itha Hajat Al-Baqara Ala Qoroniha”, its translation is ”If the cow performs Hajj on its horns” (See Table 7) was correctly understood by five out of six models: ALLaM, LLaMA, Jais, Qwen, and ACEGPT all provided the correct answer, while only Mistral failed.
+This proverb is relatively widespread, and its meaning, “something impossible to achieve”, is often documented in public sources, likely increasing its representation in model training data.
+In contrast, proverbs that are more regionally isolated or culturally specific, such as the Southern fill-in-the-blank proverb : ”Khattet fi maa wa aqbas fi heid”, its translation is ”Plan in water and pinch in a rock” (See Table 8), saw a complete failure across all six models.
+None of the models, i.e., ALLaM, LLaMA, Jais, Mistral, Qwen, and ACEGPT, selected the correct option, suggesting a consistent gap in exposure to underrepresented regional idioms, especially those originating from areas with fewer accessible textual resources.
+Table 7: Model answers to a proverb meaning question in the general dialect.
+Content Type: Proverb
+Question Type: Meaning
+Region: General
+Models’ Answer: [Uncaptioned image] ALLaM: B , [Uncaptioned image] LLaMA: B, [Uncaptioned image] Jais: B, [Uncaptioned image] Mistral: C, [Uncaptioned image] Qwen: B, [Uncaptioned image] ACEGPT: B
+Table 8: Model answers to a fill-in-the-blank proverb in the southern dialect.
+Content Type: Proverb
+Question Type: Fill-in-the-Blank
+Region: Southern
+Models’ Answer: [Uncaptioned image] ALLaM: D , [Uncaptioned image] LLaMA: B, [Uncaptioned image] Jais: D, [Uncaptioned image] Mistral: D, [Uncaptioned image] Qwen: A, [Uncaptioned image] ACEGPT: D
+Table 9: Model answers to a location recognition word question in the central dialect.
+Content Type: Word
+Question Type: Location Recognition
+Region: Central
+Models’ Answer: [Uncaptioned image] ALLaM : A , [Uncaptioned image] LLaMA : B, [Uncaptioned image] Jais : A, [Uncaptioned image] Mistral : B, [Uncaptioned image] Qwen : C, [Uncaptioned image] ACEGPT : D
 Additionally, the study revealed notable differences in the models’ ability to comprehend regional dialects, particularly in how they responded to specific tasks based on region and question type.
-For example, when tasked with identifying a location-specific term (Table 9) from the Central region, only two models, ALLaM and Jais, were able to correctly associate it with the Al-Qassim region.
+For example, when tasked with identifying a location-specific term (See Table 9) from the Central region, only two models, ALLaM and Jais, were able to correctly associate it with the Al-Qassim region.
 The other models, however, defaulted to more familiar urban areas, like Riyadh and Jeddah, likely due to the prevalence of these dialects in the training materials.
@@
 6 Discussion
-This research presents Absher, a novel benchmark aimed at evaluating LLMs on Saudi dialects—a linguistically rich yet underrepresented subset of Arabic.
-Unlike prior work such as Dallah [14], Palm [18], and AL-QASIDA [22], which aggregate dialects or focus on broader Arabic settings, Absher uniquely provides fine-grained, region-specific coverage of Saudi dialects, including Central, Weastern, Southern, Eastern, and Northern varieties.
+This research presents Absher, a novel benchmark aimed at evaluating LLMs on Saudi dialects, a linguistically rich yet underrepresented subset of Arabic.
+Unlike prior work such as Dallah [alwajih2024dallah], Palm [alwajih2025palm], and AL-QASIDA [robinson2024qasida], which aggregate dialects or focus on broader Arabic settings, Absher uniquely provides fine-grained, region-specific coverage of Saudi dialects, including Central, Western, Southern, Eastern, and Northern varieties.
+Absher’s strength lies in its foundation: over 18,000 multiple-choice questions derived from a curated dataset of dialectal words, phrases, and proverbs sourced from these five major Saudi regions.
+The benchmark systematically spans six distinct task categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
 Our findings highlight notable patterns in how LLMs handle Saudi dialects.
 Location Recognition tasks yielded the highest performance, likely due to the concrete and frequently occurring nature of geographic references in training data, which aids retrieval and identification.
-In contrast, True/False questions were the most challenging—especially in low-context settings—reflecting the inherent difficulty of binary reasoning tasks that require nuanced cultural and contextual understanding beyond surface-level lexical cues.
-Furthermore, Qwen’s strong performance—especially in Meaning and Contextual Usage questions—demonstrates its robust semantic understanding and ability to accurately interpret dialectal expressions in context.
-Across content types, tasks involving Proverbs consistently achieved higher accuracy than those based on individual Words or Phrases.
-This suggests that models benefit from the richer contextual and cultural cues provided by longer, more nuanced inputs.
-For instance, ALLAM’s relatively high performance in Cultural Interpretation of proverbs highlights its capacity for deeper socio-cultural reasoning, likely attributed to its Arabic-centric training.
+In contrast, True/False questions were the most challenging, especially in low-context settings, reflecting the inherent difficulty of binary reasoning tasks that require nuanced cultural and contextual understanding beyond surface-level lexical cues.
+Furthermore, Qwen’s strong performance, especially in Meaning, Contextual Usage, and Cultural Interpretation questions, demonstrates its robust semantic understanding and ability to accurately interpret dialectal expressions in context.
+Across content types, performance varied by model rather than being uniformly highest for any single type.
+While some models (e.g., ALLaM) achieved relatively stronger results on Proverbs, others, such as Qwen, performed best on Word-level questions, indicating that certain models benefit more from the richer contextual and cultural cues provided by longer and more nuanced inputs.
+For instance, ALLAM’s relatively high performance in Cultural Interpretation of proverbs highlights its ability for deeper socio-cultural reasoning, likely attributed to its Arabic-centric training.
 The results reveal a noticeable gap between multilingual and Arabic-native models.
@@
 Their ability to generalize across dialectal variations, despite not being trained explicitly on Saudi dialects, reflects the benefits of exposure to diverse language data.
-These results align with observations from Palm [18] and related studies, where multilingual training correlated with higher robustness on non-standard Arabic input.
+These results align with observations from Palm [alwajih2025palm] and related studies, where multilingual models correlated with higher robustness on non-standard Arabic input.
 Nevertheless, Arabic-native models demonstrated strengths in specific areas.
@@
 These patterns suggest that domain-specific pretraining provides advantages in interpreting traditional or idiomatic expressions, even if it does not generalize as effectively across diverse question types.
-Such findings align with results from the AraDiCE benchmark [3], where models exhibited improved performance when test data resembled the training distribution.
-To further demonstrate model differences, consider a Cultural Interpretation question based on the proverb:
-{RLtext}
-“اللي ما يعرف الصقر يشويه”
-“He who doesn’t recognize the falcon cooks it”
-While ALLAM accurately selected the metaphorical interpretation (“Those who lack knowledge may misuse valuable things”), other models like Jais and ACEGPT selected overly literal or generic options.
+Such findings align with results from the AraDiCE benchmark [Baly2023], where models exhibited improved performance when test data resembled the training distribution.
+To further demonstrate model differences, consider a Cultural Interpretation question based on the proverb: ”Ally ma ya’arf alsger yshwaeh”, meaning ”He who doesn’t recognize the falcon cooks it” While ALLAM accurately selected the metaphorical interpretation (“Those who lack knowledge may misuse valuable things”), other models like Jais and ACEGPT selected overly literal or generic options.
 This reflects ALLAM’s stronger grounding in socio-cultural expressions—a critical strength when applying LLMs in culturally sensitive settings.
@@
 However, neither category achieves consistently strong performance across all dialects and task types.
-These insights underscore the need for more representative and balanced datasets, as well as tailored adaptation strategies, to improve model alignment with dialectal Arabic at the national level.
-7 Ethical Considerations and Limitations
-7.1 Ethical Considerations
-This study utilizes publicly available data sources to evaluate the dialectal understanding of LLMs within the Saudi context.
-Care was taken to ensure that no private, identifiable, or sensitive information was included in the dataset.
-Given that language models may produce hallucinated or inaccurate outputs—particularly in low-resource or culturally nuanced scenarios—we stress that these outputs should not be used for critical decision-making without human oversight.
-The evaluation framework was carefully designed to minimize cultural bias and to avoid reinforcing stereotypes.
-Throughout the project, we maintained a commitment to transparency and responsible AI usage, acknowledging the inherent limitations of current LLMs when handling dialectal and culturally embedded language.
-Ultimately, our goal is to promote the development of fair, inclusive, and culturally aware language technologies.
+This discrepancy can be attributed to structural imbalances in the representation of dialectal forms within the training data, where widely documented or urban-centered dialects are more prevalent, whereas regionally specific or culturally embedded expressions remain underrepresented.
+Furthermore, performance patterns indicate that model success is influenced not only by linguistic coverage but also by the alignment between task type and the type of cultural knowledge required.
+These insights underscore the need for more representative and balanced datasets, alongside tailored adaptation strategies and controlled evaluation methods, to improve model alignment with dialectal Arabic at the national level and to better understand why certain models excel in some contexts while struggling in others.
 6.1 Limitations
@@
 That highlights the need for more diverse and evenly distributed resources to comprehensively capture all Saudi dialectal variations in future research.
-Future research should leverage more extensive computational infrastructure, experiment with prompt engineering strategies, and incorporate broader expert reviews to further enhance the accuracy, reliability, and cultural validity of the benchmark evaluations.
-8 Conclusion and Future Work
-This paper introduced Absher, a novel benchmark designed to evaluate the capabilities of large language models (LLMs) in understanding the diverse dialects and rich cultural expressions across Saudi Arabia.
-Unlike existing Arabic benchmarks that center on Modern Standard Arabic (MSA) or treat dialects at a coarse level, Absher offers fine-grained, region-specific evaluation through over 18,000 carefully constructed multiple-choice questions.
+Future research should leverage more extensive computational infrastructure, experiment with prompt engineering strategies and fine-tuning methods.
+It should also incorporate broader expert reviews to further enhance the accuracy, reliability, and cultural validity of the benchmark evaluations.
+6.2 Ethical Considerations
+This study utilizes publicly available data sources to evaluate the dialectal understanding of LLMs within the Saudi context.
+Care was taken to ensure that no private, identifiable, or sensitive information was included in the dataset.
+To guarantee this, all dialectal content—including words, phrases, and proverbs— was sourced exclusively from the publicly accessible Moajam website.
+This resource functions as a digital repository of aggregated dialectal vocabulary and cultural expressions.
+We explicitly excluded any conversational data, social media content, or personally generated user text to safeguard against the unintentional inclusion of private information or the leakage of toxic language that is often prevalent in online datasets.
+The primary ethical risk in dialectal evaluation is the potential for linguistic discrimination, where underrepresented dialects are implicitly or explicitly associated with lower social status or performance failure by the model.
+To mitigate this critical risk, we implemented two concrete measures.
+First, the Absher dataset was rigorously designed to ensure fair coverage across the five major Saudi regions (Central, Western, Southern, Eastern, and Northern), deliberately challenging the common bias in NLP datasets that over-represent urban or dominant dialects.
+Second, the evaluation tasks focus strictly on testing linguistic competence (meaning, contextual usage, and cultural interpretation) rather than social or demographic inference, preventing the model from being scored on tasks that could encourage or propagate negative stereotypes associated with any specific regional dialect.
+Given that language models may produce hallucinated or inaccurate outputs—particularly in low-resource or culturally nuanced scenarios—we stress that these outputs should not be used for critical decision-making without human oversight.
+The evaluation framework was carefully designed to minimize cultural bias and to avoid reinforcing stereotypes.
+Finally, our commitment to responsible AI usage extended to our human validation processes.
+The validation, which was performed by native speakers from each of the five regions to confirm dialectal authenticity, was governed.
+Throughout the project, we maintained a commitment to transparency and responsible AI usage, acknowledging the inherent limitations of current LLMs, with the ultimate goal of promoting the development of fair, inclusive, and culturally aware language technologies.
+6.3 Conclusion and Future Work
+This paper introduced Absher, a novel benchmark designed to evaluate the capabilities of LLMs in understanding the diverse dialects and rich cultural expressions across Saudi Arabia.
+Unlike existing Arabic benchmarks that focus on Modern Standard Arabic (MSA) or treat dialects at a coarse level, Absher offers fine-grained, region-specific evaluation through over 18,000 carefully constructed multiple-choice questions.
 These questions span six task types—ranging from meaning identification to cultural interpretation—and are grounded in authentic dialectal content from various Saudi regions.
+These questions are derived from a curated dataset of dialectal words, phrases, and proverbs sourced from five major Saudi regions, and systematically span six distinct task categories: Meaning, True/False, Fill-in-the-Blank, Contextual Usage, Cultural Interpretation, and Location Recognition.
+This comprehensive design enables multifaceted evaluation across both linguistic and cultural dimensions.
 Our evaluation of several state-of-the-art LLMs revealed significant variation in performance across dialects, question types, and linguistic units.
@@
 Looking forward, future work will expand Absher to include additional dialects, question formats, and model types.
-We also plan to explore few-shot and instruction-tuned settings to assess model adaptability in low-resource scenarios.
+We plan to conduct more in-depth evaluations using alternative strategies, such as few-shot and fine-tuning scenarios, to better explore models’ capabilities across dialectal variations and cultural contexts.
+In addition, we aim to perform deeper causal analyses to understand the factors driving model performance across dialects and task types, providing clearer insights into how linguistic and cultural features influence LLM behavior.
 Integrating sociolinguistic features and region-specific datasets may further improve alignment with real-world language use.
@@
 We provided several illustrative examples from our benchmark in this paper, and the complete dataset will be made publicly available for research purposes upon acceptance.
+CRediT authorship contribution statement
+Renad Al-Monef: Methodology, Data Curation, Software, Validation, Formal analysis, Visualization, Writing - original draft.
+Hassan Alhuzali: Conceptualization, Supervision, Validation, Writing - review & editing.
+Nora Alturayeif: Supervision, Writing - review & editing.
+Ashwag Alasmari: Conceptualization, Supervision, Validation, Writing - review & editing.
 Declaration of competing interest
@@
 After using this tool/service, we reviewed and edited the content as needed and take full responsibility for the content of the publication.
+The authors would like also to extend their appreciation to the Deanship of Research and Graduate Studies at King Khalid University for funding this work through small group research under grant number RGP1/69/46.
+The authors also express their sincere gratitude to Al Bandari Ali Alqahtani , Zainab Abdulaziz Dalim for their invaluable support and insightful contributions during the evaluation of Absher benchmark
 Appendices
@@
 Appendix B Annotation Guideline
-Figure 8: The interface designed using Google Sheets, which was shared with annotators to evaluate the questions.
-Appendix C Detailed Performance
+Appendix C The Annotation interface
+Appendix D Detailed Performance
+Table 10: Detailed evaluation of the results across content and question types, with the best results highlighted in dark green to represent the highest scores.
 Model Content Type Question Type Accuracy Precision Recall F1-score
@@
 Location Recognition 38.67% 30.48% 37.80% 28.22%
-Table 10: Evaluation Results of the Model’s Performance